find a point on the x axis that is equidistant from (-1,-1) and (7,7)
Answer by stanbon(75887)

find a point on the x axis that is equidistant from (-1,-1) and (7,7)
----
Let the point be (x,0); solve for "x":
-----
Equation:
distance = distance
(x+1)^2 + (-1)^2 = (x-7)^2 + 7^2
---
x^2 + 2x + 2 = x^2 - 14x + 98
16x = 96
x = 6
---
Point: (6,0)
====================
Cheers,
Stan H.